OpenStudy (yrivers36):

How do I find the domain? f(x)=3/x=6

OpenStudy (anonymous):

Are you sure you typed the function correctly? At the moment you have the solution right there, if you're supposed to find the domain you'd typically have a function like this: f(x) = 1/x + 2x Without actually having a answer like you have: f(x) = 3/x = 6, since that just gives x = 1/2.

OpenStudy (yrivers36):

sorry hit wrong key f(x)=3/x+6

OpenStudy (anonymous):

To find the domain you want to know what value x can have while the function still is valid. So the only thing you got to worry about is 3/x. When does 3/x not evaluate to a correct value?

OpenStudy (yrivers36):

I am not understanding this problem

OpenStudy (anonymous):

It's not as hard as it seems, the solution is that the function is valid for all values of x except for x = 0, that would evaluate to 3/0 which is invalid. That's basically how you're supposed to see it, what can x be without making the function break the rules of math. :)

OpenStudy (yrivers36):

infinite?

OpenStudy (anonymous):

Try look at this, more of an explanation of domains and ranges. http://mathforum.org/library/drmath/sets/select/dm_domain_range.html

OpenStudy (yrivers36):

Thank you